OpenStudy (anonymous):

giving out medal to anyone who answers these. really need help 1.Name the subset(s) of real numbers to which each number belongs: 1a. squareroot of 43 1b. -14 1c. 0 2.order the numbers from least to greatest: -1/6, -0.3, squareroot of 1, -2/13, 7/8 3. for the following problems name the property illistruated. 3a. 9+1.3=1.3+9 3b. -1*6=-6 3c. a*0=0 3d. 12+(9+1)=(12+9)+1 3e. 5*(3*2)=(5*3)*2 3f. 4+0=4 3g. 6*1=6 4. using the formula r=d/t, where d is distance in miles, r is rate, t is time in hrs, at which rate must you travel to cover 252mi in 4.5hrs? explain.

OpenStudy (skullpatrol):

That's alot of questions.

OpenStudy (skullpatrol):

You don't know how to do any of them?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

even if someone could help me with a few of them thatd be awesome, id appreciate it a lot.. but no ones really helping me so /:

OpenStudy (anonymous):

"4. using the formula r=d/t, where d is distance in miles, r is rate, t is time in hrs, at which rate must you travel to cover 252mi in 4.5hrs? explain. so, d = 252 t = 4.5 then r = 252/4.5

OpenStudy (anonymous):

thankyou so much can yu help me with any of the others?

OpenStudy (anonymous):

"2.order the numbers from least to greatest: -1/6, -0.3, squareroot of 1, -2/13, 7/8 put the fractions into decimals, then order the fractions based on the values of the decimals -1/6 = -0.166 -0.3 = -0.3 squareroot of 1 = 1 -2/13 = -0.1538 7/8 = .875

OpenStudy (anonymous):

so, 0.1538, .875, 0.166, 0.3, and 1 ?

OpenStudy (anonymous):

don't forget the negative signs, try again tho

OpenStudy (anonymous):

is .875 less than .166? noh, so you gotta swap those around

OpenStudy (anonymous):

i dont know :(

OpenStudy (anonymous):

okay, i think i got it thanks

OpenStudy (anonymous):

ur welcome